Air at 14.7 psia, 40 oF enters an ideal four-stroke Diesel cycle with a compression ratio of 20 and a displacement of 300 in3. It runs at 2100 RPM and the maximum temperature is 3300 oF. The fuel has an energy content of 19,360 Btu/lbm.

a. The power output of the engine in Hp b. The cycle thermal efficiency c. The fuel consumption rate in lbm/hr
Answer #1

ANSORY Given that the ay P214--1p31Q 0-101352Mpa T1 -40°F 277596K. MK=26-) V220 N=200rpm Displacement. = 300 in3 20.0049173 TTH = 2088-706 - T : 2088706 (8.8101704 T4 = 874-7488 Qy=CvCT4-T) = 6-718 874-748–27715%) = 498-756 kJ /Kg. thermal efficiencyV -V = 45-675 (0-786–0-039) =998-226&pa. Break Powe3 = MCP XA XLXN . 2x60 = 998-226x0.00441 x2100 120 - 85-712 klo =115012 hp